# Swin UNETR++: Advancing Transformer-Based Dense Dose Prediction Towards Fully Automated Radiation Oncology Treatments

### v2

@@ Abstract
Radiation Dense Dose Prediction, Vision Transformer, Swin UNETR, Swin UNETR++, Dual Cross-Attention
@@ 1 Introduction
Radiotherapy remains a critical tool in the fight against cancer, employing sophisticated techniques like volumetric-modulated arc therapy (VMAT) for precise radiation delivery. Yet, even as artificial intelligence (AI) has emerged to assist with organ segmentation in the clinic, the development of patient-specific treatment plans remains a highly complicated and time-consuming task that lies outside the reach of current AI-based approaches. This complexity originates from trade-off decisions that must balance the maximum radiation dose to the tumor while minimizing the dose to surrounding normal tissues. Furthermore, converting physician treatment plan intent to deliverable linear accelerator or “robot” sequenced motions involves solving a complex inverse optimization problem using proprietary simulated annealing algorithms within dedicated treatment planning systems. To address this problem, several groups have worked to develop techniques to translate anatomical information, i.e., 3D images and associated organ segmentation, to a predicted radiation dose distribution that matches clinical constraints (Liu et al., 2021a; Nguyen et al., 2019c). Competitions such as Open Knowledge-Based Planning Grand Challenge (OpenKBP) (Babier et al., 2021) have been initiated to challenge the community to make progress on this problem, but current approaches still depend on fully convolutional models that may not be sufficient for generating deliverable treatment plans due to the lack of large receptive fields. The goal of this work is to transition fully convolutional 3D dose prediction to attention-based.
We introduced Swin UNETR++, a novel deep learning architecture that aims to supplant the extant fully convolutional neural network (FCNN) approaches in 3D dose prediction. This new model is backboned by Swin UNETR (Hatamizadeh et al., 2021) and incorporates a lightweight 3D Dual Cross-Attention (DCA) module designed to bridge the semantic gap between encoder and decoder features in the skip connection. This architecture excels at capturing both channel and spatial dependencies, thereby providing a more holistic view of the patient’s anatomy for accurate dose prediction. Swin UNETR++ was rigorously evaluated using the OpenKBP Grand Challenge dataset and trained using a hybrid of mean squared error loss $\mathcal{L}_{\text{MSE}}$ and a novel global differentiable approximation of the dose-volume histogram loss $\mathcal{L}_{\text{DVH\_global}}$. In terms of performance metrics, Swin UNETR++ shows near-state-of-the-art results, outpacing most existing models in both the validation and test phases, and its implementation marks a significant advancement in automating the radiotherapy planning process. This study not only outlines the architecture and performance of Swin UNETR++ but also opens avenues for further research in incorporating more nuanced attention mechanisms into treatment planning.
@@ 2 Related Work
After the OpenKBP Grand Challenge dataset from (Babier et al., 2021) was released, several teams have shown the strength of full CNN. Nguyen et al. (2019b) proposed Hierarchically Densely Connected U-net, HD U-net, based on U-net (Ronneberger et al., 2015) and DenseNet (Huang et al., 2017). Soomro et al. (2021) introduced DeepDoseNet, which integrates Resnet (He et al., 2016) and dilated densenet blocks in their architecture. Liu et al. (2021a) proposed cascade 3D model, C3D, which ranked No.1 in OpenKBP Grand Challenge on both $\overline{S_{\text{DVH}}}$ and $\overline{S_{\text{Dose}}}$. In C3D, two U-nets are connected in series, and the second U-net refines the coarse prediction made by the first U-net. As ViT (Dosovitskiy et al., 2020) and Swin UNETR (Hatamizadeh et al., 2021) result in state-of-the-art performance in classification and segmentation tasks, we use Swin UNETR++ to demonstrate the superiority of the transformer-based approach in 3D radiation dose estimation.
@@ 3.1 Dataset
The OpenKBP Grand Challenge, dataset from (Babier et al., 2021), contains 340 patients treated by intensity-modulated radiation therapy (IMRT) for head-and-neck cancer. Each patient was prescribed a combination of three possible radiation doses: 70 Gy to the gross disease planning target volume (PTV70), 63 Gy to the intermediate-risk planning target volumes (PTV63), and 56 Gy to the elective planning target volumes (PTV56), with organs-at-risk (OARs) masks including the brain stem, spinal cord, right parotid, left parotid, esophagus, larynx, and mandible. In addition, CT, the ground-truth 3D radiation dose distribution, and possible dose mask, which serves as a body mask defining the area that can receive dose, are provided for each patient. All image data are downsampled to 128 × 128 × 128 voxel with voxel dimensions around 3.5 mm × 3.5 mm × 2 mm. In terms of dataset distribution, OpenKBP Grand Challenge assigned patients No.1–200 for training, No.201–240 for validation, and No.241–340 for testing.
@@ 3.2 Model Architecture
Figure 2: DCA module internal structure
fig:arch2
@@ 3.2.1 Swin UNETR
Swin UNETR has recently emerged as a promising method for medical image segmentation challenges, registering state-of-the-art performance in specialized domains such as glioma delineation (Hatamizadeh et al., 2021), Medical Segmentation Decathlon (Simpson et al., 2019), and Beyond The Cranial Vault Segmentation Challenge (Tang et al., 2022). The model combines the computational advantages of the Swin Transformer (Liu et al., 2021b), known for its shifted window attention mechanisms, with the domain-specific robustness of UNETR (Hatamizadeh et al., 2022), a specialized variant of U-Net for 3D volumetric medical imaging tasks. This combination benefits from spatial hierarchy representation but also adheres to computational efficiency guidelines. However, despite these merits, our analysis has identified an overlooked shortcoming—namely, the semantic gap between the encoder and decoder features propagated through the skip connections. The discrepancies between low-level features and high-level semantics in U-Net-based architectures are corroborated by (Pang et al., 2019; Wang et al., 2022). Resolving this semantic inconsistency could improve the accuracy of 3D radiation dose prediction, which requires more details and clinical understanding than semantic segmentation.
@@ 3.2.2 Swin UNETR++
[…]
@@
  • 3D Patch Embedding. We used the same patch size defined by (Ates et al., 2023). Given $n$ encoder stages and encoder features $E_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times\frac{D}{2^{i-1}}}$ where $i\in\{1,2,...,n\}$, the flattened patches for stage $i$ are extracted by 3D $1\times 1\times 1$ depth-wise convolutions and 3D average pooling layer as following:
@@
By combining the 3D DCA module with the Swin UNETR architecture in the skip connection, we proposed Swin UNETER++, as illustrated in Figure LABEL:fig:arch1 and Figure LABEL:fig:arch2. It provides an improvement in medical image analysis and is specifically tailored for dense radiation dose prediction. It is also the first time the 3D DCA module was applied outside semantic segmentation.
The 3D DCA module has clinical importance as well. CCA refines the model’s understanding of the channel-wise relationship between different clinically significant volumes, similar to the prior work of overlap volume histogram (OVH) that quantified the distance between different OARs and PTVs (Wu et al., 2009). SCA establishes the close-and-distant relationships within each clinically significant volume. Together, CCA and SCA aim to heighten treatment efficacy while minimizing radiation-induced complications, contributing to a safer, more effective therapeutic regimen.
@@ 3.3.1 Data Preprocessing and Augmentation
Inspired by (Liu et al., 2021a), we first clipped the CT images into a fixed range between -1024 and 1500 HU and then normalized them by 1000 HU during preprocessing. Any missing clinically significant volume masks for each patient are represented by a 3D tensor with all zeros during data loading.
During training, we applied data augmentation with MONAI (Medical Open Network for Artificial Intelligence) by Cardoso et al. (2022) to increase our model’s generalizability and avoid overfitting in the following order: (1)Random axis flip which randomly flips the spatial axis of CT, structured masks, and possible dose mask (probability=0.5). (2)Random affine which performs random shearing in a range of $(0.1,0.1,0.1)$ with zero-padding and nearest-neighbor interpolation on CT, structured masks, and possible dose mask (probability=0.2). (3)Random zoom scales CT, structured masks, and possible dose mask within a range of $[0.9,1.3]$ with nearest-neighbor interpolation (probability=0.5). (4)Random Guassian smooth is applied only to CT (probability=0.2). (5). Random contrast is applied only to CT (probability=0.2).
@@ 3.3.2 Loss Function
The DVH loss $\mathcal{L}_{\text{DVH}}$, as shown in (Nguyen et al., 2020), is a loss function that takes into account a common clinical tool for assessing the dose to each organ in a treatment plan, the dose volume histogram (DVH). As the traditional DVH calculation is not differentiable, a differentiable approximate dose-volume histogram $\widetilde{\text{DVH}}$ is used. Given a set of clinically significant volume masks $M_{v}$, where $v\in V$ and $V=\text{PTV}\cup\text{OAR}$, with the number of $M_{v}$ each patient has $n_{v}$ and the number of bins $n_{t}$ in $\widetilde{\text{DVH}}$, we can obtain:
@@
We introduced a variant of the DVH loss, the global DVH loss $\mathcal{L}_{\text{DVH\_global}}$. The global DVH condenses the integral dose across the patient, i.e., the possible dose mask segmentation, to a single line on the DVH plot. The loss then considers the difference of this integral dose distribution between the predicted and ground truth distributions. This loss was shown to have value as it considers the radiation dose across the entire patient’s body as opposed to individual clinically segmented volumes that are sparse and account for only a fraction of the total patient volume. Given the possible dose mask $M_{d}$, $\mathcal{L}_{\text{DVH\_global}}$ is computed as following:
@@ 3.3.3 Hyperparameter Setting
Adam optimizer (Kingma and Ba, 2014) was initialized with a learning rate of 0.0002 and beta values of (0.5, 0.999). To adaptively adjust the learning rate during the training process, we utilized the Cosine Annealing with Warm Restarts scheduler. The initial restart period was set to 100 epochs, with a multiplicative factor of 1 and a minimum learning rate of 0.00002. Furthermore, we fixed seed to 42, thereby ensuring deterministic behavior for stochastic operations and facilitating experiment reproducibility.
@@ 4.1.1 Dose Score
Provided in the OpenKBP Grand Challenge by (Babier et al., 2021), the Dose Score $\overline{S_{\text{Dose}}}$ is the mean absolute difference between the predicted radiation dose distribution $D_{\text{pred}}$ and the ground-truth radiation dose distribution $D_{\text{gt}}$ based on the possible dose mask $M_{\text{d}}$ among all validation or test patients $P_{t}$ where $t=1,2,...,n$.
@@
$=\frac{\sum_{i,j,k}\left|D_{\text{pred}}(i,j,k)-D_{\text{gt}}(i,j,k)\right|M_{\text{d}}(i,j,k)}{\sum_{i,j,k}M_{\text{d}}(i,j,k)}$ (17)
@@ 4.1.2 DVH Score
DVH Score $\overline{S_{\text{DVH}}}$ is designed to measure the quality of the predicted radiation dose distribution clinically. According to (Babier et al., 2021), the radiation dose distribution for each OAR is evaluated by:
@@
Finally, the mean absolute difference between the predictions and the ground truth radiation dose distributions is computed based on these metrics for each patient in the validation and the test dataset with their existing OAR and PTV masks.
@@ 4.2 Qualitative Metrics
To evaluate the clinical reliability of the model’s predictions, we defined the absolute difference between the ground-truth and predicted dose distribution within 3 Gy as clinically acceptable and the absolute difference outside this range as clinically unacceptable based on (Sher et al., 2021; Mashayekhi et al., 2021).
@@ 4.2.1 Volume-wise Clinical Acceptance
We consider the predicted radiation dose distribution for a volume to be volume-wise clinically acceptable VA if and only if the absolute dose distribution difference between the predicted and the ground-truth within the given volume is less than 3 Gy. The function VA() has two possible outputs: 1 represents volume-wise clinically acceptable while 0 represents volume-wise clinically unacceptable. Given a set of validation or test patients $P$ and a set of all clinically significant volumes $V$, the volume-wise clinical acceptance rate $R_{\text{VA}}$ for a volume $v\in V$ is defined as:
@@ 4.2.2 Patient-wise Clinical Acceptance
Clinically, medical dosimetrists try to achieve Pareto optimal plans given individual patient’s unique anatomy and physicians’ prioritization, but different prioritization can lead to different “optimal” plans (Nguyen et al., 2019a; Kyroudi et al., 2020). To simulate this real-world plan variations, we first find the absolute difference between the predicted and ground-truth radiation dose distribution for all clinically significant volumes $V$ each patient has and then average those absolute differences. We consider a patient to be patient-wise clinically acceptable PA if and only if the average difference is less than 3 Gy. The function PA() has two possible outputs: 1 represents patient-wise clinically acceptable while 0 represents patient-wise clinically unacceptable. Given a set of all validation or test patients $P$, the average patient-wise clinical acceptance rate $\overline{R_{\text{PA}}}$ for an entire validation or test dataset is defined as:

### v3

@@ Abstract
Radiation Dense Dose Prediction, Vision Transformer, Swin UNETR++, Dual Cross-Attention
@@ 1 Introduction
Radiotherapy remains a critical tool in the fight against cancer, employing sophisticated techniques like intensity-modulated radiation therapy (IMRT) for precise radiation delivery. Yet, even as artificial intelligence (AI) has emerged to assist with organ segmentation in the clinic, the development of patient-specific treatment plans remains a highly complicated and time-consuming task that lies outside the reach of current AI-based approaches. This complexity originates from trade-off decisions that must balance the maximum radiation dose to the tumor while minimizing the dose to surrounding normal tissues. Furthermore, converting physician treatment plan intent to deliverable linear accelerator or “robot” sequenced motions involves solving a complex inverse optimization problem using proprietary simulated annealing algorithms within dedicated treatment planning systems. To address this problem, several groups have worked to develop techniques to translate anatomical information, i.e., 3D images and associated organ segmentation, to a predicted radiation dose distribution that matches clinical constraints (Liu et al., 2021a; Nguyen et al., 2019c). Competitions such as Open Knowledge-Based Planning Grand Challenge (OpenKBP) (Babier et al., 2021) have been initiated to challenge the community to make progress on this problem, but current approaches still depend on fully convolutional models that may not be sufficient for generating deliverable treatment plans due to the lack of large receptive fields. The goal of this work is to transition fully convolutional 3D dose prediction to attention-based.
We introduced Swin UNETR++, a novel deep learning architecture that aims to supplant the extant fully convolutional neural network (FCNN) approaches in 3D dose prediction. This new model is backboned by Swin UNETR (Hatamizadeh et al., 2021) and incorporates a lightweight 3D Dual Cross-Attention (DCA) module designed to bridge the semantic gap between encoder and decoder features in the skip connection. This architecture excels at capturing both channel and spatial dependencies, thereby providing a more holistic view of the patient’s anatomy for accurate dose prediction. Swin UNETR++ was rigorously evaluated using the OpenKBP dataset and trained using a hybrid of mean squared error loss $\mathcal{L}_{\text{MSE}}$ and a novel global differentiable approximation of the dose-volume histogram loss $\mathcal{L}_{\text{DVH\_global}}$. In terms of performance metrics, Swin UNETR++ shows near-state-of-the-art results, outpacing most existing models in both the validation and test phases, and its implementation marks a significant advancement in automating the radiotherapy planning process. This study not only outlines the architecture and performance of Swin UNETR++ but also opens avenues for further research in incorporating more nuanced attention mechanisms into treatment planning.
@@ 2 Related Work
After the OpenKBP dataset was released, several teams have shown the strength of full CNN. Nguyen et al. (2019b) proposed Hierarchically Densely Connected U-net (HD U-net). Soomro et al. (2021) introduced DeepDoseNet, which integrates Resnet (He et al., 2016) and dilated densenet blocks in their architecture. Liu et al. (2021a) proposed cascade 3D model, C3D, which ranked No.1 on OpenKBP leaderboard for both $\overline{S_{\text{DVH}}}$ and $\overline{S_{\text{Dose}}}$. In C3D, two U-nets are connected in series, and the second U-net refines the coarse prediction made by the first U-net. As ViT (Dosovitskiy et al., 2020) and Swin UNETR result in state-of-the-art performance in classification and segmentation tasks, we use Swin UNETR++ to demonstrate the superiority of the transformer-based approach in 3D radiation dose estimation.
@@ 3.1 Dataset
The OpenKBP dataset contains 340 patients treated by IMRT for head-and-neck cancer. Each patient was prescribed a combination of three possible radiation doses: 70 Gy to the gross disease planning target volume (PTV70), 63 Gy to the intermediate-risk planning target volumes (PTV63), and 56 Gy to the elective planning target volumes (PTV56), with organs-at-risk (OARs) masks including the brain stem, spinal cord, right parotid, left parotid, esophagus, larynx, and mandible. In addition, CT, the ground-truth 3D radiation dose distribution, and a possible dose mask contouring the body are provided for each patient. All image data are downsampled to 128 × 128 × 128 voxels with voxel dimensions around 3.5 mm × 3.5 mm × 2 mm. In terms of dataset distribution, OpenKBP Grand Challenge assigned patients No.1–200 for training, No.201–240 for validation, and No.241–340 for testing.
@@ 3.2 Model Architecture
[…]
fig:arch2 Refer to caption
@@ 3.2.1 Swin UNETR
Swin UNETR has recently emerged as a promising method for medical image segmentation challenges, registering state-of-the-art performance in specialized domains such as glioma delineation, Medical Segmentation Decathlon (Simpson et al., 2019), and Beyond The Cranial Vault Segmentation Challenge (Tang et al., 2022). The model combines the computational advantages of the Swin Transformer (Liu et al., 2021b), known for its shifted window attention mechanisms with the domain-specific robustness of UNETR (Hatamizadeh et al., 2022). This combination benefits from spatial hierarchy representation but also adheres to computational efficiency guidelines. However, despite these merits, our analysis has identified an overlooked shortcoming—namely, the semantic gap between the encoder and decoder features propagated through the skip connections. The discrepancies between low-level features and high-level semantics in U-Net-based architectures are corroborated by Pang et al. (2019) and Wang et al. (2022). Resolving this semantic inconsistency could improve the accuracy of 3D radiation dose prediction, which requires more details and clinical understanding than semantic segmentation.
@@ 3.2.2 Swin UNETR++
Research has been looking at using attention mechanisms to improve the performance of U-Net-based architectures. For instance, Mou et al. (2021) introduced CS2-Net, which utilized spatial and channel attention to facilitate the segmentation of curvilinear structures. The DANet (Fu et al., 2019) integrated position and channel attention module in scene segmentation. Recently, Ates et al. (2023) proposed a light-weight 2D plug-and-play Dual Cross-Attention (DCA) module which has demonstrated its efficiency in reducing the semantic gap among common fully convolutional U-Net-based architecture.
@@
3D Patch Embedding. We used the same patch size defined by Ates et al. (2023). Given $n$ encoder stages and encoder features $E_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times\frac{D}{2^{i-1}}}$ where $i\in\{1,2,...,n\}$, the flattened patches for stage $i$ are extracted by 3D $1\times 1\times 1$ depth-wise convolutions and 3D average pooling layer as following:
@@
By combining the 3D DCA module in Figure LABEL:fig:arch2 with the Swin UNETR architecture in the skip connection, we proposed Swin UNETER++ illustrated in Figure LABEL:fig:arch1. It improves medical image analysis and is specifically tailored for dense radiation dose prediction. It is also the first time the 3D DCA module was applied outside semantic segmentation.
The 3D DCA module has clinical importance as well. CCA refines the model’s understanding of the channel-wise relationship between different clinically significant volumes, similar to the prior work of overlap volume histogram (Wu et al., 2009) that quantified the distance between different OARs and PTVs. SCA establishes the close-and-distant relationships within each clinically significant volume. Together, CCA and SCA aim to heighten treatment efficacy while minimizing radiation-induced complications, contributing to a safer, more effective therapeutic regimen.
@@ 3.3.1 Data Preprocessing and Augmentation
Inspired by Liu et al. (2021a), we first clipped the CT images into a fixed range between -1024 and 1500 HU and then normalized them by 1000 HU during preprocessing. Any missing volume masks for each patient are represented by a 3D zero tensor.
During training, we applied data augmentation with MONAI (Cardoso et al., 2022) to increase our model’s generalizability and avoid overfitting in the following order: (1)Random axis flip which randomly flips the spatial axis of CT, structured masks, and possible dose mask (probability=0.5). (2)Random affine which performs random shearing in a range of $(0.1,0.1,0.1)$ with zero-padding and nearest-neighbor interpolation on CT, structured masks, and possible dose mask (probability=0.2). (3)Random zoom scales CT, structured masks, and possible dose mask within a range of $[0.9,1.3]$ with nearest-neighbor interpolation (probability=0.5). (4)Random Guassian smooth is applied only to CT (probability=0.2). (5)Random contrast is applied only to CT (probability=0.2).
@@ 3.3.2 Loss Function
The DVH loss $\mathcal{L}_{\text{DVH}}$ (Nguyen et al., 2020) is a loss function that takes into account a common clinical tool for assessing the dose to each organ in a treatment plan, the dose volume histogram (DVH). As the traditional DVH calculation is not differentiable, a differentiable approximate dose-volume histogram $\widetilde{\text{DVH}}$ is used. Given a set of $n_{v}$ clinically significant volume masks $M_{v}$, where $v\in V=\text{PTV}\cup\text{OAR}$, and the number of bins $n_{t}$, we can obtain:
@@
We introduced a variant of the DVH loss, the global DVH loss $\mathcal{L}_{\text{DVH\_global}}$. The global DVH condenses the integral dose across the patient, i.e., the possible dose mask segmentation $M_{d}$, to a single line on the DVH plot. The loss then considers the difference of this integral dose distribution between the predicted and ground truth distributions. This loss considers the radiation dose across the entire patient’s body as opposed to individual clinically segmented volumes that are sparse and account for only a fraction of the total patient volume. $\mathcal{L}_{\text{DVH\_global}}$ is computed as follows:
@@ 3.3.3 Hyperparameter Setting
Adam optimizer (Kingma and Ba, 2014) was initialized with a learning rate of 2e-4 and beta values of (0.5, 0.999). To adaptively adjust the learning rate during the training process, we utilized the Cosine Annealing with Warm Restarts scheduler. The initial restart period was set to 100 epochs, with a multiplicative factor of 1 and a minimum learning rate of 2e-5. Furthermore, we fixed the seed to 42, thereby ensuring deterministic behavior for stochastic operations and facilitating experiment reproducibility.
@@ 4.1.1 Dose Score
Provided by OpenKBP, the Dose Score $\overline{S_{\text{Dose}}}$ is the mean absolute difference between the predicted radiation dose distribution $D_{\text{pred}}$ and the ground-truth radiation dose distribution $D_{\text{gt}}$ based on the possible dose mask $M_{\text{d}}$ for each patient $P_{t}$ in either validation or test datasets.
@@
$=\frac{\sum_{i,j,k}\left|D_{\text{pred}}(i,j,k)-D_{\text{gt}}(i,j,k)\right|M_{\text{d}}(i,j,k)}{\sum_{i,j,k}M_{\text{d}}(i,j,k)}$
@@ 4.1.2 DVH Score
DVH Score $\overline{S_{\text{DVH}}}$ (Babier et al., 2021) is designed to measure the quality of the predicted radiation dose distribution clinically. The radiation dose distribution for each OAR is evaluated by:
@@
Finally, the mean absolute difference between the predictions and the ground truth radiation dose distributions is computed based on the above metrics for each patient using their existing OAR and PTV masks.
@@ 4.2 Qualitative Metrics
To evaluate the clinical reliability of the model’s predictions, we defined the absolute difference between the ground-truth and predicted dose distribution within 3 Gy as clinically acceptable and the absolute difference outside this range as clinically unacceptable (Sher et al., 2021; Mashayekhi et al., 2021).
@@ 4.2.1 Volume-wise Clinical Acceptance
We consider the predicted radiation dose distribution for a volume to be volume-wise clinically acceptable VA if and only if the absolute dose distribution difference between the predicted and the ground-truth within the given volume is less than 3 Gy. The function $\text{VA}()$ has two possible outputs: 1 represents volume-wise clinically acceptable while 0 represents unacceptable. Given a set of validation or test patients $P$ and a set of clinically significant volumes $V$, the volume-wise clinical acceptance rate $R_{\text{VA}}$ for a volume $v\in V$ is defined as:
@@ 4.2.2 Patient-wise Clinical Acceptance
Clinically, medical dosimetrists try to achieve Pareto optimal plans given individual patient’s unique anatomy and physicians’ prioritization, but different prioritization can lead to different “optimal” plans (Nguyen et al., 2019a; Kyroudi et al., 2020). To simulate these real-world plan variations, we first find the absolute difference between the predicted and ground-truth radiation dose distribution for all clinically significant volumes $V$ of each patient $p$ and then average those absolute differences. We consider a patient to be patient-wise clinically acceptable PA if and only if the average difference is less than 3 Gy. The function PA() has two possible outputs: 1 represents patient-wise clinically acceptable while 0 represents patient-wise clinically unacceptable. Given a set of all validation or test patients $P$, the average patient-wise clinical acceptance rate $\overline{R_{\text{PA}}}$ for an entire validation or test dataset is defined as: